# Budgeted Online Continual Learning by Adaptive Layer Freezing and Frequency-based Sampling

### v1

@@ Abstract
The majority of online continual learning (CL) advocates single-epoch training and imposes restrictions on the size of replay memory. However, single-epoch training would incur a different amount of computations per CL algorithm, and the additional storage cost to store logit or model in addition to replay memory is largely ignored in calculating the storage budget. Arguing different computational and storage budgets hinder fair comparison among CL algorithms in practice, we propose to use floating point operations (FLOPs) and total memory size in Byte as a metric for computational and memory budgets, respectively, to compare and develop CL algorithms in the same ‘total resource budget.’ To improve a CL method in a limited total budget, we propose adaptive layer freezing that does not update the layers for less informative batches to reduce computational costs with a negligible loss of accuracy. In addition, we propose a memory retrieval method that allows the model to learn the same amount of knowledge as using random retrieval in fewer iterations. Empirical validations on the CIFAR-10/100, CLEAR-10/100, and ImageNet-1K datasets demonstrate that the proposed approach outperforms the state-of-the-art methods within the same total budget. : Equal contribution. $\leavevmode\nobreak\ \leavevmode\nobreak\ {}^{\dagger}$: Corresponding author.
@@ 1 Introduction
For the practicality of online continual learning (CL), the scenarios assumed by most online CL methods impose resource restrictions, such as the single training epoch and the size of the replay memory that limits the number of streamed samples that can be stored (Koh et al., 2022; Wang et al., 2022a). Although the ‘one-epoch training’ may give a rough sense of the computational constraint, the actual computation budget to train the models differs per method (Prabhu et al., 2023; Verwimp et al., 2023; Ghunaim et al., 2023) since each method requires a different amount of computations in a single epoch. Several rehearsal-based CL methods require additional storage to store the previous models and logits (i.e., the unnormalized model output vector) (Buzzega et al., 2020; Koh et al., 2023; Zhou et al., 2023), which was usually not included in the memory budget, which mainly considers the size of episodic memory to store samples in previous tasks. To this end, we compare CL methods with the same computational and memory budget considering all storage and computational costs. We argue that the total budget of memory and computation will ensure the practicality of the proposed online CL algorithms.
For a fair comparison in computational budget, we use training FLOPs per sample as the metric instead of the number of epochs, since some methods require a lot of computations in a single epoch, while others do not. FLOPs is an exact computational budget regardless of implementation details (Korthikanti et al., 2023), following (Zhao et al., 2023; Ghunaim et al., 2023).
@@
For the same memory budget, we use an aggregated budget for various forms of extra storage, including replay memory, logits, and model parameters. Following (Zhou et al., 2023), we convert all additional storage costs into byte and sum them up to obtain the actual memory cost. Upon comparing the results under fair cost, we found that the performances were often different from what was reported in the original papers proposing each method.
@@
Figure 1: Comparison of CL methods w/o total constraint (left) and w/ total constraint (right) on CIFAR-10 Gaussian setup. In the left plot, we compare CL methods with the same number of iterations and the same episodic memory size, i.e., conventional setup. In the right plot, we compare CL methods with the same training FLOPs per sample and a fixed storage budget that includes both episodic memory and model storage, i.e., our total-constrained CL setup. Compared to the conventional setup, aL-SAR shows improved performance under the total-constrained setup, since it can utilize the saved computational cost for additional model training.
@@
Taking into account the total memory and computational budget, we additionally propose an online CL method using a computation-aware layer freezing strategy which only requires an additional cost of 0.02% compare to the replay memory size in ImageNet). Specifically, we propose to selectively learn (or freeze) layers per a mini-batch based on the amount of information already learned in the previous mini batches, to reduce computations. In particular, as the number of layers being frozen increases, training computation can be saved at the cost of losing information for the mini-batch since frozen layers cannot acquire any information. Considering the trade-off between computation and information, we propose ‘adaptive layer freezing’, which chooses the best layers to freeze by maximizing the Fisher Information (FI) gained by the model for each batch, given a fixed computation budget. Unlike previous freezing methods (Lee et al., 2019; Hayes et al., 2020; Yuan et al., 2022), which predefine the rule of when and which layers to freeze, not generalizing to various datasets and models, we determine the appropriate layers to freeze in each forward pass considering the varying information contained in each batch.
@@
For the loss of accuracy by the reduced computational cost, we propose a novel sample retrieval method to improve accuracy with negligible training cost. While several online CL methods, such as MIR (Aljundi et al., 2019a) and ASER (Shim et al., 2021), aim to retrieve more informative training batches than random retrieval from replay memory, these methods require forward/backward passes of a large candidate set to select informative batches. This results in significantly higher computational overhead, with MIR and ASER requiring 4 times and 3 times more computation than ER, respectively. The application of these methods would nullify the effort of reducing the computation by the proposed adaptive layer freezing method. To this end, we propose to retrieve samples that the model has not learned much about information stored in episodic memory for training without incurring additional computational costs. To quantify the degree of learning, we employ the frequency of recent usage of each sample in training and the similarity of the gradients between classes. They are acquired during the training process, i.e., no need for additional inference.
@@
  • Proposing a computationally efficient adaptive layer freezing that maximizes Fisher Information per computation, as well as a memory retrieval strategy that promotes the retrieval of samples that the model has not learned much.
@@ 4.2 Quantitative Analysis
We evaluated CL methods, including aL-SAR, with strictly restricted computation and memory budgets as specified in Sec. 4.1. Note that while we set the training iterations of aL-SAR to be the same as other baselines, aL-SAR adaptively freezes layers, resulting in fewer FLOPs consumed. For experiments on various computational and memory budgets, we use the relatively small CIFAR datasets to cover a wide range of given budgets. To validate our methods on large datasets and datasets with temporal domain shift, we also show experiments on ImageNet and CLEAR datasets. Note that these experiments are conducted on various CL setups, including Gaussian task setup, Disjoint task setup, and domain-incremental setup.
@@
To investigate the scalability of CL methods on large datasets and temporal domain shifts, we report the performances on ImageNet-1K and CLEAR-10/100 with a fixed computational and memory budget in Tab. 1. In the ImageNet-1K Gaussian setup, aL-SAR outperforms other methods with slightly fewer FLOPs, and in the ImageNet-1K Disjoint setup, aL-SAR shows similar performance with other methods while using significantly fewer FLOPs. In this setup, training distribution changes continuously, so the model has to constantly adapt to the new distribution, resulting in less freezing and more benefit from the fast adaptation enabled by the retrieval method. In contrast, in the ImageNet-1K Disjoint setup, the distribution shift only occurs at the task boundaries, so fast adaptation by retrieval has a smaller effect, but aL-SAR freezes more layers after each task has been sufficiently learned.
@@
Gaussian Setup Disjoint Setup
$A_{\text{AUC}}\ \uparrow$ $A_{\text{last}}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{\text{last}}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{\text{last}}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{\text{last}}\ \uparrow$
EWC (Kirkpatrick et al., 2017) $20.05$ $32.55$ $44.17$ $36.11$ 62.40$\pm$0.17 71.25$\pm$1.29 29.62$\pm$0.44 40.36$\pm$1.34
ER (Rolnick et al., 2019) $20.01$ $32.60$ $44.13$ $36.85$ 64.00$\pm$0.06 72.28$\pm$1.15 31.56$\pm$0.42 44.90$\pm$0.74
ER-MIR (Aljundi et al., 2019a) $7.45$ $17.04$ $35.48$ $35.83$ 62.29$\pm$0.17 70.52$\pm$1.13 23.24$\pm$0.14 34.92$\pm$1.16
REMIND (Hayes et al., 2020) $15.58$ $22.42$ $37.42$ $33.14$ 62.64$\pm$0.04 70.68$\pm$0.90 26.76$\pm$0.30 30.99$\pm$0.55
DER++ (Buzzega et al., 2020) $23.51$ $29.82$ $40.30$ $37.09$ 61.30$\pm$0.77 71.52$\pm$1.02 25.83$\pm$0.43 39.89$\pm$0.89
OCS (Yoon et al., 2022) $6.90$ $19.57$ $27.50$ $30.41$ 58.69$\pm$0.25 68.69$\pm$1.21 15.24$\pm$1.27 22.52$\pm$2.02
LiDER (Bonicelli et al., 2022) $27.65$ $34.66$ $32.68$ $31.73$ 58.36$\pm$0.75 69.16$\pm$1.21 28.75$\pm$0.28 39.92$\pm$1.29
X-DER (Boschini et al., 2023) $27.46$ $34.47$ $32.34$ $31.41$ 58.19$\pm$0.23 69.70$\pm$0.97 23.01$\pm$0.15 33.70$\pm$0.70
MEMO (Zhou et al., 2023) $11.15$ $19.98$ $39.30$ $33.35$ 60.29$\pm$0.31 68.69$\pm$1.21 20.33$\pm$0.28 25.53$\pm$0.92
CAMA (Kim et al., 2024) $7.94$ $17.44$ $35.38$ $36.02$ 60.97$\pm$0.52 71.66$\pm$1.18 26.27$\pm$0.37 39.42$\pm$0.67
CCL-DC (Wang et al., 2024) $19.70$ $32.12$ $44.07$ $35.98$ 61.08$\pm$0.55 71.25$\pm$0.82 23.81$\pm$0.32 35.07$\pm$0.67
aL-SAR (Ours) 33.94 38.02 45.08 37.25 67.61$\pm$0.45 74.36$\pm$0.80 38.65$\pm$0.30 47.56$\pm$0.37
Table 1: Quantitative comparison between different CL methods on CIL setup. aL-SAR outperforms baselines with a lower computational budget and the same storage budget. Specifically, while all other baselines consume 112,000, 716, and 5,250 TFLOPs, aL-SAR consumes 93,000, 598, and 4,059 TFLOPs in the ImageNet-1K, CLEAR-10, and CLEAR-100, respectively. The memory budget is fixed at 5,736MB, 564MB, and 1,148MB in ImageNet-1K, CLEAR-10, and CLEAR-100, respectively.
@@
Methods Bongard-HOI Bongard-OpenWorld
@@ 5 Conclusion
We take reproducibility in deep learning very seriously and highlight some of the contents of the manuscript that might help to reproduce our work. We will definitely release our implementation of the proposed method in Sec. 3, the data splits and the baselines used in our experiments in Sec. 4.
@@ A.11 Experimental Results on Tasks from Various Datasets
A.12 Details on Training Multi-Modal Large Language Models
@@
Beyond the class-incremental learning (CIL) setup, we extend aL to the multi-modal concept-incremental task (MCIL). Specifically, we evaluate the Bongard-HOI (Jiang et al., 2022) and Bongard-OpenWorld (Wu et al., 2024) benchmarks, which feature multiple ’concepts’ (e.g., ride a bike, The top of a snow-covered mountain) in each benchmark. We split the concepts into 5 disjoint tasks for the MCIL setup. These benchmarks leverage two key attributes of traditional Bongard problems (Depeweg et al., 2018): (1) the capacity for few-shot concept learning and (2) reasoning that is dependent on context. The former entails the ability to derive visual concepts from a limited number of examples, while the latter suggests that the classification of a query image can differ based on the context provided (i.e., the positive and negative support sets). In the Bongard problem, with a positive support set and a negative support set, we tackle two specific tasks: (1) Which concept is exlusively represented by the positive support set? and (2) For a given query image, does it belong to the positive or negative support set? We denote these tasks as CA (Concept Answering) and P/N, respectively.
@@
In the P/N task, the answers are either ‘Positive’ or ‘Negative’, so we utilize accuracy as the evaluation metric. Conversely, for the CA tasks, we compare the ground truth sentences with the output sentences generated by the MLLM model. To assess this, we employ CIDEr (Vedantam et al., 2015), which evaluates how closely a predicted sentence aligns with a set of ground-truth sentences. Similar to class-incremental tasks, we also measure the area under the curve (AUC) performance and the final performance within this multi-modal framework.
@@
For the MCIL setup, we set the episodic memory size to 500 and use the Adam optimizer with a learning rate of $5\times 10^{-5}$ and a Constant LR scheduler. The number of images in the support set is set to 2 and 3 images for each positive and negative support set in Bongard-HOI and Bongard-OpenWorld, respectively. We use the LLaVA-1.5-7B model and train it on NVIDIA A100 80GB GPUs.
@@
No freezing $26.43\pm 1.32$ $\textbf{23.86}\pm\textbf{1.49}$ 1478.78
aL (Ours) $\textbf{27.48}\pm\textbf{2.90}$ $23.76\pm 1.79$ 1308.72 (-11.5%)
@@
In Tab. 3 in Sec. 4.2, we empirically demonstrate the effectiveness of aL in the Bongard-HOI-P/N and Bongard-OpenWorld-P/N tasks. Here, we apply aL to the Bongard-OpenWorld-CA task and summarize the results in Tab. 9. Similar to Bongard-HOI-P/N and Bonagd-OpenWorld-P/N tasks, aL effectively reduces training costs. Since our proposed adaptive layer freezing method is independent of downstream tasks, model architecture, and task identity information, we expect it can be applied to a variety of tasks and architectures.
A.13 Comparison between aL and Other Freezing Methods

### v2

@@ Abstract
The majority of online continual learning (CL) advocates single-epoch training and imposes restrictions on the size of replay memory. However, single-epoch training would incur a different amount of computations per CL algorithm, and the additional storage cost to store logit or model in addition to replay memory is largely ignored in calculating the storage budget. Arguing different computational and storage budgets hinder fair comparison among CL algorithms in practice, we propose to use floating point operations (FLOPs) and total memory size in Byte as a metric for computational and memory budgets, respectively, to compare and develop CL algorithms in the same ‘total resource budget.’ To improve a CL method in a limited total budget, we propose adaptive layer freezing that does not update the layers for less informative batches to reduce computational costs with a negligible loss of accuracy. In addition, we propose a memory retrieval method that allows the model to learn the same amount of knowledge as using random retrieval in fewer iterations. Empirical validations on the CIFAR-10/100, CLEAR-10/100, and ImageNet-1K datasets demonstrate that the proposed approach outperforms the state-of-the-art methods within the same total budget. Furthermore, we validate its effectiveness in the Multi-modal Concept incremental Learning setup with multimodal large language models, such as LLaVA-1.5-7B. Code is available at https://github.com/snumprlab/budgeted-cl.
@@ 1 Introduction
In a realistic scenario for continual learning (CL), data arrive in a streaming manner, prompting interest in online CL, which assumes one or a few samples arrive at a time. To effectively learn from new data while mitigating catastrophic forgetting (McCloskey & Cohen, 1989) of previously learned knowledge, various CL methods have been proposed, including replay-based approaches (Bang et al., 2021; Seo et al., 2024), network expansion methods (Wu et al., 2022; Zhou et al., 2023), and distillation-based methods (Koh et al., 2023; Wang et al., 2024).
@@
For the practicality of online CL, most online CL methods impose resource restrictions, such as the single training epoch and limited replay memory, which restrict the number of streamed samples stored (Koh et al., 2022; Wang et al., 2022a). While the ‘one-epoch training’ may give a rough sense of the computational constraint, the actual budget varies across methods (Prabhu et al., 2023; Ghunaim et al., 2023) since each method requires a different amount of computations in a single epoch. Several rehearsal-based CL methods require additional storage to store the previous models and logits (Buzzega et al., 2020; Zhou et al., 2023), which was usually not included in the memory budget, which mainly considers the size of episodic memory to store samples in previous tasks. To this end, we compare CL methods with the same computational and memory budget considering all storage and computational costs. We argue that the total budget of memory and computation will ensure the practicality of the proposed online CL algorithms.
@@
For a fair comparison in computational budget, we use training FLOPs per sample instead of the number of epochs, as some methods require significantly more computations per epoch than others. FLOPs provide an exact measure of computational budget regardless of implementation details (Korthikanti et al., 2023), following Zhao et al. (2023); Ghunaim et al. (2023). For the same memory budget, we use an aggregated budget for various forms of extra storage, including replay memory, logits, and model parameters, converting them into bytes to obtain the actual memory cost, following  Zhou et al. (2023). Upon comparing the results under fair cost, we found that the performances were often different from what was reported in the original papers proposing each method.
@@
Figure 1: Comparison of CL methods w/o total constraint (left) and w/ total constraint (right) on CIFAR-10 Gaussian setup. In the left plot, we compare CL methods with the same number of iterations and the same episodic memory size, i.e., conventional setup. In the right plot, we compare CL methods with the same training FLOPs and a fixed storage budget that includes both episodic memory and model storage, i.e., our total-constrained CL setup. Compared to the conventional setup, aL-SAR shows improved performance under the total-constrained setup, since it can utilize the saved computational cost for further training. $A_{\text{AUC}}$ and $A_{\text{LAST}}$ refer to the area under the curve of accuracy and last accuracy (i.e., accuracy at the end of CL), respectively.
@@
Considering the total memory and computational budget, we propose an online CL method using a computation-aware layer freezing strategy. Specifically, we propose to selectively learn (or freeze) layers per a mini-batch based on the previously learned information to reduce computation In particular, as more layers are frozen, training computation can be saved at the cost of losing information for the mini-batch since frozen layers cannot acquire any information. Considering this trade-off, we propose ‘adaptive layer freezing’, which chooses the best layers to freeze by maximizing the Fisher Information (FI) gained by the model for each batch, given a fixed computation budget. Unlike previous freezing methods (Lee et al., 2019; Hayes et al., 2020; Yuan et al., 2022), which predefine the rule of when and which layers to freeze, not generalizing to various datasets and models, we adaptively select layers to freeze based on the varying information in each batch.
@@
For the loss of accuracy by the reduced computational cost, we propose a novel sample retrieval method to improve accuracy with negligible training cost. While several online CL methods, such as MIR (Aljundi et al., 2019a) and ASER (Shim et al., 2021), aim to retrieve more informative training batches than random retrieval from replay memory, they require forward/backward passes of a large candidate set to select informative batches, increasing computation by 4 times and 3times compared to ER (Rolnick et al., 2019), respectively. Applying these methods negates the computational savings of the proposed adaptive layer freezing approach. To this end, we propose to retrieve samples that the model has not learned much about information stored in episodic memory for training without incurring additional computational costs. To quantify the degree of learning, we employ the frequency of recent usage of each sample in training and the similarity of the gradients between classes. They are acquired during the training process, i.e., no need for additional inference.
@@
  • Proposing a computationally efficient adaptive layer freezing that maximizes FI per computation, as well as a memory retrieval strategy that prioritizes samples that the model has learned least.
@@ 4.2 Quantitative Analysis
We evaluated CL methods, including aL-SAR, with strictly restricted computation and memory budgets as specified in Sec. 4.1. Note that while we set the training iterations of aL-SAR to be the same as other baselines, aL-SAR adaptively freezes layers, resulting in fewer FLOPs consumed. For experiments on various computational and memory budgets, we use the relatively small CIFAR datasets to cover a wide range of given budgets. To validate our methods on large datasets and datasets with temporal domain shift, we also show experiments on ImageNet and CLEAR datasets. Note that these experiments are conducted on various CL setups, including Gaussian task setup, Disjoint task setup, and domain-incremental setup. Furthermore, we applied our adaptive layer freezing method to LLaVA-v1.5-7B, demonstrating a significant reduction in the computational cost of training the multi-modal large language model while preserving performance.
@@
To investigate the scalability of CL methods on large datasets and temporal domain shifts, we report the performances on ImageNet-1K Gaussian setup and CLEAR-10/100 with a fixed computational and memory budget in Tab. 1. In the ImageNet-1K Gaussian setup, aL-SAR outperforms other methods with slightly fewer FLOPs. In this setup, training distribution shifts continuously, so the model has to constantly adapt to the new distribution, resulting in less freezing and more benefit from the fast adaptation enabled by the retrieval method.
@@
$A_{\text{AUC}}\ \uparrow$ $A_{\text{last}}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{\text{last}}\ \uparrow$ $A_{\text{AUC}}\ \uparrow$ $A_{\text{last}}\ \uparrow$
EWC (Kirkpatrick et al., 2017) $20.05$ $32.55$ 62.40$\pm$0.17 71.25$\pm$1.29 29.62$\pm$0.44 40.36$\pm$1.34
ER (Rolnick et al., 2019) $20.01$ $32.60$ 64.00$\pm$0.06 72.28$\pm$1.15 31.56$\pm$0.42 44.90$\pm$0.74
ER-MIR (Aljundi et al., 2019a) $7.45$ $17.04$ 62.29$\pm$0.17 70.52$\pm$1.13 23.24$\pm$0.14 34.92$\pm$1.16
REMIND (Hayes et al., 2020) $15.58$ $22.42$ 62.64$\pm$0.04 70.68$\pm$0.90 26.76$\pm$0.30 30.99$\pm$0.55
DER++ (Buzzega et al., 2020) $23.51$ $29.82$ 61.30$\pm$0.77 71.52$\pm$1.02 25.83$\pm$0.43 39.89$\pm$0.89
OCS (Yoon et al., 2022) $6.90$ $19.57$ 58.69$\pm$0.25 68.69$\pm$1.21 15.24$\pm$1.27 22.52$\pm$2.02
LiDER (Bonicelli et al., 2022) $27.65$ $34.66$ 58.36$\pm$0.75 69.16$\pm$1.21 28.75$\pm$0.28 39.92$\pm$1.29
X-DER (Boschini et al., 2023) $27.46$ $34.47$ 58.19$\pm$0.23 69.70$\pm$0.97 23.01$\pm$0.15 33.70$\pm$0.70
MEMO (Zhou et al., 2023) $11.15$ $19.98$ 60.29$\pm$0.31 68.69$\pm$1.21 20.33$\pm$0.28 25.53$\pm$0.92
CAMA (Kim et al., 2024) $7.94$ $17.44$ 60.97$\pm$0.52 71.66$\pm$1.18 26.27$\pm$0.37 39.42$\pm$0.67
CCL-DC (Wang et al., 2024) $19.70$ $32.12$ 61.08$\pm$0.55 71.25$\pm$0.82 23.81$\pm$0.32 35.07$\pm$0.67
aL-SAR (Ours) 33.24 35.02 67.61$\pm$0.45 74.36$\pm$0.80 38.65$\pm$0.30 47.56$\pm$0.37
Table 1: Quantitative comparison between different CL methods on CIL setup. aL-SAR outperforms baselines with a lower computational budget and the same storage budget. Specifically, while all other baselines consume 112,000, 716, and 5,250 TFLOPs, aL-SAR consumes 93,000, 598, and 4,059 TFLOPs in the ImageNet-1K Gaussian setup, CLEAR-10, and CLEAR-100, respectively. The memory budget is fixed at 5,736MB, 564MB, and 1,148MB in ImageNet-1K, CLEAR-10, and CLEAR-100, respectively.
@@
Methods Bongard-HOI-P/N Bongard-OpenWorld-P/N
@@ 5 Conclusion
To further facilitate the reproduction, we provide open-source implementations of our proposed method (Sec.3), along with data splits and baseline models used in our experiments (Sec.4), available at https://github.com/snumprlab/budgeted-cl.
Acknowledgment
This work was partly supported by AI Center, Samsung Electronics, and the IITP grants (No.RS-2022-II220077, No.RS-2022-II220113, No.RS-2022-II220959, No.RS-2021-II211343 (SNU AI), No.RS-2021-II212068 (AI Innov. Hub)) funded by the Korea government (MSIT).
@@ A.11 Experimental Results on Tasks from Various Datasets
A.12 Comparison with SOTA Efficient CL Methods
(a) CIFAR-10 Gaussian task setup                                      (b) CIFAR-10 Disjoint task setup
(a) CIFAR-100 Gaussian task setup                                      (b) CIFAR-100 Disjoint task setup
A.13 Details on Training Multi-Modal Large Language Models
@@
Beyond the class-incremental learning (CIL) setup, we extend aL to the multi-modal concept-incremental learning (MCIL) setup. Specifically, we evaluate the Bongard-HOI (Jiang et al., 2022) and Bongard-OpenWorld (Wu et al., 2024) benchmarks, which feature multiple ’concepts’ (e.g., ride a bike, The top of a snow-covered mountain) in each benchmark. We split the concepts into 5 disjoint tasks for the MCIL setup. These benchmarks leverage two key attributes of traditional Bongard problems (Depeweg et al., 2018): (1) the capacity for few-shot concept learning and (2) reasoning that is dependent on context. The former entails the ability to derive visual concepts from a limited number of examples, while the latter suggests that the classification of a query image can differ based on the context provided (i.e., the positive and negative support sets). In the Bongard problem, with a positive support set and a negative support set, we tackle two specific tasks: (1) Which concept is exlusively represented by the positive support set? and (2) For a given query image, does it belong to the positive or negative support set? We denote these tasks as CA (Concept Answering) and P/N, respectively.
@@
For the Positive/Negative (P/N) task, where answers are either Positive or Negative, we use accuracy as the evaluation metric. Similarly, for the Bongard-HOI CA task, which has simple text answers (e.g., kick a ball), we also use accuracy as the metric. In contrast, for the Bongard-OpenWorld-CA task, where answers are full sentences, we compare the generated sentences from the MLLM model with ground truth sentences. Specifically, following  Liu et al. (2023); Yu et al. (2024), we use GPT-3.5 to assess response quality, evaluating helpfulness, relevance, accuracy, and detail. It provides an overall score from 1 to 10, with higher scores indicating better performance. We provide details on the prompts used for GPT-4 evaluation in Sec. A.14. Similar to class-incremental tasks, we also measure the area under the curve (AUC) performance and the final performance within this multi-modal framework.
@@
For the MCIL setup, we assume infinite episodic memory size and use the Adam optimizer with a learning rate of $5\times 10^{-5}$ and a Constant LR scheduler. The number of images in the support set is set to 2 for each positive and negative support set in Bongard-HOI and Bongard-OpenWorld. We use the LLaVA-1.5-7B model and train it on NVIDIA A100 80GB GPUs.
Methods Bongard-HOI-CA Bongard-OpenWorld-CA
@@
No freezing 74.46$\pm$2.77 72.41$\pm$2.02 788.35 73.72$\pm$0.49 73.26$\pm$0.56 1478.78
aL (Ours) 75.27$\pm$2.15 72.31$\pm$3.26 715.43 (-9.25%) 73.06$\pm$1.15 73.92$\pm$2.51 1401.28 (-5.24%)
@@
In Tab. 3 in Sec. 4.2, we empirically demonstrate the effectiveness of aL in the Bongard-HOI-P/N and Bongard-OpenWorld-P/N tasks.
Effect of aL in CA tasks.
In addition to the P/N tasks presented in Table 3, we also apply aL to the CA tasks and summarize the results in Table 9. Similar to the P/N tasks, aL effectively reduces training costs in CA tasks. However, compared to the relatively simple P/N tasks, which involve binary classification, the CA tasks require generating text or sentence-level answers, making them more complex. As a result, aL leads to a relatively smaller reduction in FLOPs in CA tasks, compared to P/N tasks. It shows that aL adaptively freezes fewer layers for more complex tasks that require more training.
Notably, our proposed adaptive layer freezing method operates independently of downstream tasks, model architecture, and task identity information, thus we expect it to be applicable across a wide range of tasks and model architectures.
A.14 Details of Prompts used for GPT-3.5 Evaluation
We use GPT-3.5 to assess the quality of predicted sentences against ground truth sentences using the following prompts:
[Ground truth]: ‘Ground-Truth sentence’ [Prediction]: ‘Prediction sentence’ Compare the ground truth and prediction from AI models, to give a correctness score for the prediction. The user asks the question about observing an image or multiple images. The images replace <image> tokens in [Input]. Please rate the relevance and accuracy of [Assistant] compared to [Ground-Truth]. The assistant receives an overall score on a scale of 1 (totally wrong) to 10 (totally right), where a higher score indicates better overall performance. Please first provide a comprehensive explanation of your evaluation explanation. In the subsequent line, please provide a single line containing the score for Assistant, in the format of Score: <an integer value on a scale 1 to 10>
A.15 Comparison between aL and Other Freezing Methods